# Vid2Coach: Transforming How-To Videos into Task Assistants

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 2.2. Skimming and Navigating How-To Video
When following how-to videos, people engage in complex navigation behaviors – rewatching certain parts to identify details and skipping irrelevant parts (Chang et al., 2021). To facilitate navigation, prior research explored segmenting based on tool usage (Truong et al., 2021; Grossman et al., 2010), intermediate results (Kim et al., 2014), and other learners’ interaction data (Yang et al., 2022). These navigation units are often visualized as multi-modal summaries with keyframes and transcript snippets  (Truong et al., 2021; Pavel et al., 2014; Chang et al., 2021), or graph-based visualization (Liu et al., 2018; Kim et al., 2023). Researchers have also explored how to contextually present these how-to video segments based on tool context (Fraser et al., 2019; Banovic et al., 2012; Nguyen and Liu, 2015) and physical space (Thoravi Kumaravel et al., 2019; Sadprasid et al., 2024). Compared to manually scrubbing through the video timeline, these navigational units make it easier to jump to meaningful moments. However, they often rely on visual interaction, which can be difficult when viewers’ eyes and hands are occupied with the physical task. Thus, researchers proposed voice-based video navigation as an alternative (Chang et al., 2021; Truong et al., 2021; Zhao et al., 2022). Building on this prior work, we segment videos into structured steps and support voice-based control. While prior systems take explicit user commands for navigation, we detect users’ step completion and proactively suggests moving to the next step – enabling mixed-initiative guidance for time-sensitive tasks like cooking.
@@ 2.3. Physical Task Assistance
People following procedural tasks like cooking often rely on voice assistants (e.g., Amazon Alexa, Google Home) to read out instructions, answer questions about ingredients, or set timers (Hwang et al., 2023; Jaber et al., 2024). With advances in speech recognition and natural language understanding, these assistants now go beyond fixed commands and can infer user intent from free-form speech. Beyond voice interactions, wearable assistants can enable richer context understanding with visual, audio, and motion sensors. For example, head-mounted AR systems can be used to disambiguate user questions through visual context (Lee et al., 2024b) and to convey the affordances of kitchen tools to low-vision users (Lee et al., 2024a). To support task assistance through a more common consumer device, Arakawa et al. developed smart watch systems that deliver context-aware support during physical tasks (Arakawa et al., 2024b, a).
@@ 2.4. Computer Vision for Accessibility
The capability to assist with physical tasks is closely linked to computer vision innovations. How-to videos have served as a valuable data source (Miech et al., 2019) for training video understanding models capable of tasks like action recognition (Ashutosh et al., 2023; Zhou et al., 2023), step localization (Xue et al., 2024b; Dvornik et al., 2023), skill assessment (Ashutosh et al., 2024a) and video question answering (Zhong et al., 2022). The recent advent of large language models has catalyzed the development of vision-language Models (VLMs) (Li et al., 2024c; Wang et al., 2024; Achiam et al., 2023; Team et al., 2023), which offer a unified natural language interface for diverse vision tasks and present a promising avenue for analyzing user videos to provide real-time assistance. Despite their potential, current VLMs face substantial hurdles when used for physical task guidance, especially for BLV users. One primary issue is hallucination (Bai et al., 2024), where models produce information not grounded in the visual evidence. Also, these models often generate guidance presupposing visual capabilities, leading to suggestions unsuitable for BLV users (Bigham et al., 2010; Song et al., 2024). Further, existing VLMs often lack fine-grained analysis capabilities (Xue et al., 2024a), struggling to monitor detailed execution progress within tasks, which is vital for useful step-by-step assistance. In this work, we carefully design task support grounded in BLV people’s needs in real-world scenarios and propose a pipeline to generate accessible task guidance.
@@ 3.1. Method
Figure 3. Timelines of utterances annotation of observational study sessions. Colors represent the type of utterances. VRT2 proactively provided more supplementary information and progress descriptions to BLV2 who was new to baking.
@@ 3.3. Study Findings
{mdframed}
[ linewidth=0.8pt, backgroundcolor=gray!10, linecolor=black, innerrightmargin=6pt, innerleftmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt ] Takeaway: Provide detailed instructions based on both the narration and visuals from the video. Structure rich information into clear, concise segments to support easier understanding.
@@
{mdframed}
[ linewidth=0.8pt, backgroundcolor=gray!10, linecolor=black, innerrightmargin=6pt, innerleftmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt ] Takeaway: Supplement instructions with tips and accessible workarounds for both immediate guidance and building reusable strategies for future tasks.
@@
{mdframed}
[ linewidth=0.8pt, backgroundcolor=gray!10, linecolor=black, innerrightmargin=6pt, innerleftmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt ] Takeaway: Provide proactive visual feedback on user progress and success, and prompt users when adjustments are needed to keep relevant actions in view.
@@
{mdframed}
[ linewidth=0.8pt, backgroundcolor=gray!10, linecolor=black, innerrightmargin=6pt, innerleftmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt ] Takeaway: Elicit users to leverage non-visual sensory cues to evaluate the progress, to build confidence and support independent task completion.
@@
{mdframed}
[ linewidth=0.8pt, backgroundcolor=gray!10, linecolor=black, innerrightmargin=6pt, innerleftmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt ] Takeaway: Address users’ diverse questions by grounding responses in both the user’s current task progress and the how-to video knowledge, enabling flexible task execution.
@@
{mdframed}
[ linewidth=0.8pt, backgroundcolor=gray!10, linecolor=black, innerrightmargin=6pt, innerleftmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt ] Takeaway: Adapt guidance to user skills, preferences, and context by adjusting the level of detail, rephrasing based on familiarity, and referencing available tools or spatial layout.
@@ 5.2. Supplementing Accessible Strategies
Retrieval Augmented Generation. General-purpose AI suggestions often assume dominant user groups, and may generate irrelevant suggestions that make BLV users feel misfit (e.g., Check if the butter looks golden brown(Alharbi et al., 2024). Even when prompted to provide accessible tips, these models reflect biased assumptions and suggest that users seek help from sighted people or offer overly generic advice without practical value – reminding them to be careful while cooking. To address these limitations, we use a multi-modal retrieval augmented generation (RAG) approach (Lewis et al., 2020) to ground AI suggestions in real-world, disability-aware resources. From the curated dataset, our pipeline first extracts text chunks to generate text embeddings, and identifies images and describes them with a VLM (Gemini 1.5 pro) to generate embeddings. At inference time, we input queries to generate step-specific workarounds, and the query also incorporates user-specific context to further personalize suggestions (e.g., vision level, cooking experience, available tools and kitchen setup) (see § 8 for details). We compute the cosine similarity between the query’s text embedding and the stored embeddings and retrieve top-3 text chunks to generate answers with the context.
@@ 5.3. Progress Monitoring
From the how-to video, our pipeline generates a set of criteria (irrelevant, in-progerss, complete, mistake) for each action and uses these to monitor the user’s progress (Figure 5). These criteria are derived using a VLM (Achiam et al., 2023) (details in §C). Instead of directly comparing frames from the how-to and user streams, we use abstracted criteria that are more robust to variations in tools, occlusions, and user behavior. As users perform each action, Vid2Coach uses these criteria to adapt its feedback: it pauses during unrelated or impromptu actions (e.g., washing hands), provides progress updates while a step is in-progress, and proactively suggests moving on once the step appears complete.
To make this system effective in real-time streamed video, we propose a 3-way categorization of actions – punctual, iterative, or durative – based on its temporal characteristics. Punctual actions (e.g.,“add 1 cup of flour”) happen quickly and may be missed in frame sampling or batched inference. For punctual actions, in-progress feedback is often untimely or unnecessary, so Vid2Coach only confirms completion. For iterative actions (e.g., “place three scoops of cookie dough”) that involve repetition, it counts visible repetitions for tracking progress. For durative actions (e.g., “cook until golden brown”) that involve gradual visual change, Vid2Coach offers real-time updates on its progression and completion. For actions where visual cues are minimal (e.g., “wait for the mixture to cool down”), Vid2Coach asks users to decide when to proceed.
To balance the accuracy and speed of realtime feedback, we adopt a dual-model approach. Inspired by WorldScribe (Chang et al., 2024), which leveraged different models with varying latency and granularity, our setup combines 1) a batch model (Gemini 2.0) for deeper reasoning with 2) a lightweight streaming model (Gemini 2.0-Live) for immediate responsiveness. The batch VLM runs every 5 seconds and looks at the five most recent frames (sampled at 1 fps) to determine user status based on pre-generated completion criteria. Unlike batch models, streaming models begin generating output token-by-token before fully processing the input (Xiao et al., 2023), making them less suited for complex reasoning but valuable for continuous monitoring. Thus, we use the streaming VLM to provide real-time, low-latency descriptions of user actions when a durative action is in progress. Users can toggle this progress feedback on or off using a voice command. We reset the model’s input context for each step to avoid memory build-up and maintain consistent inference speed.
@@ 5.4. Intent Elicitation & Question Answering
In addition to automated feedback, Vid2Coach also responds to users’ open-ended, voice-based queries. To ensure responsiveness, we use the streaming model’s voice activity detection (VAD) to immediately interrupt any ongoing generation and prioritize the user’s spoken query. Inspired by patterns observed in our formative study (§3), we implement a few-shot classification prompt that identifies the user’s intent from both explicit utterances (e.g., “Does this look complete?”) and implicit expressions (e.g., “I can’t tell which of these is sugar”). We classify user utterances into the following five intent types and trigger a different function call to handle each request:
@@ 6.1. Generating Instructions
Results. Table 1 summarizes the evaluation results. Overall, Vid2Coach outperformed baselines by identifying more task-relevant atomic facts and omitting fewer details present in the original narration. As illustrated in Figure 6, Vid2Coach surfaced concrete visual details that can support BLV users to accurately follow the steps (e.g., describing the use of a measuring spoon, amount of bacon used, and how they’re presented). These details are especially important for BLV users, who might otherwise miss critical visual cues and inaccurately follow the action.
@@ 7.4. Limitations and Opportunities
Future Opportunities. Participants also saw Vid2Coach’s potential to enable independent task skill learning in everyday lives. Compared to human assistants, participants found Vid2Coach more readily available, task-relevant, and nonjudgmental. Several participants emphasized the personal motivation and sense of achievement they felt by completing tasks on their own, even when human help was accessible. P4 noted “Even if sighted volunteers are available 24/7 for free, I want to know I can achieve it on my own.” P11, who saw the potential in personalization of Vid2Coach said “Unlike human assistants whose description styles change every call, I can train it [Vid2Coach] to be personalized to me.” P5 also reflected on societal opportunities “I think it can help blind students get more jobs. A lot of sighted people assume blind people can’t learn new skills quickly or follow manuals. But with this, we can.”
@@ 9. Discussion
9.1. Videos as a Knowledge Source
@@ 9.3. Supporting New Domains
While Vid2Coach uses curated accessibility guidelines and BLV-shared practices for common tasks like cooking and crafting, resources for less common tasks (e.g., furniture assembly, planting) are limited. Future systems can incorporate community-sourced knowledge, while also actively eliciting and supporting the sharing of such knowledge. This could involve creating lightweight platforms for BLV users to exchange non-visual workarounds, similar to how blind communities share tips through forums and mailing lists. Peer-to-peer knowledge exchange may surface effective strategies that may not yet be captured in formal accessibility resources.
9.4. Beyond Voice Assistant

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: The 38th Annual ACM Symposium on User Interface Software and Technology; September 28-October 1, 2025; Busan, Republic of Koreabooktitle: The 38th Annual ACM Symposium on User Interface Software and Technology (UIST ’25), September 28-October 1, 2025, Busan, Republic of Koreadoi: 10.1145/3746059.3747612isbn: 979-8-4007-2037-6/2025/09isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Accessibility
@@ 2.2. Skimming and Navigating How-To Video
When following how-to videos, people engage in complex navigation behaviors – rewatching certain parts to identify details and skipping irrelevant parts (Chang et al., 2021). To facilitate navigation, prior research explored segmenting based on tool usage (Truong et al., 2021; Grossman et al., 2010), intermediate results (Kim et al., 2014), and other learners’ interaction data (Yang et al., 2022). These navigation units are often visualized as multi-modal summaries with keyframes and transcript snippets  (Truong et al., 2021; Pavel et al., 2014; Chang et al., 2021), or graph-based visualization (Liu et al., 2018; Kim et al., 2023). Researchers have also explored how to contextually present these how-to video segments based on tool context (Fraser et al., 2019; Banovic et al., 2012; Nguyen and Liu, 2015) and physical space (Thoravi Kumaravel et al., 2019; Sadprasid et al., 2024). Compared to manually scrubbing through the video timeline, these navigational units make it easier to jump to meaningful moments. However, they often rely on visual interaction, which can be difficult when viewers’ eyes and hands are occupied with the physical task. Thus, researchers proposed voice-based video navigation (Chang et al., 2021; Truong et al., 2021; Zhao et al., 2022). Building on this prior work, we segment videos into structured steps and support voice-based control. While prior systems take explicit user commands for navigation, we detect users’ step completion and proactively suggests moving to the next step – enabling mixed-initiative guidance for time-sensitive tasks like cooking.
@@ 2.3. Physical Task Assistance
People following procedural tasks like cooking often rely on voice assistants (e.g., Amazon Alexa, Google Home) to read out instructions (Ju et al., 2001), answer questions about ingredients, or set timers (Hwang et al., 2023; Jaber et al., 2024). With advances in speech recognition and natural language understanding, these assistants now go beyond fixed commands and can infer user intent from free-form speech. Beyond voice interactions, wearable assistants can enable richer context understanding with visual, audio, and motion sensors (Chi et al., 2007). For example, head-mounted AR systems can be used to disambiguate user questions through visual context (Lee et al., 2024b) and to convey the affordances of kitchen tools to low-vision users (Lee et al., 2024a). To support task assistance through a more common consumer device, Arakawa et al. developed smart watch systems that deliver context-aware support during physical tasks (Arakawa et al., 2024b, a).
@@ 2.4. Computer Vision for Accessibility
The capability to assist with physical tasks is closely linked to computer vision innovations. How-to videos have served as a valuable data source (Miech et al., 2019) for training video understanding models capable of tasks like action recognition (Ashutosh et al., 2023; Zhou et al., 2023), step localization (Xue et al., 2024b; Dvornik et al., 2023), skill assessment (Ashutosh et al., 2024a) and video question answering (Zhong et al., 2022). The recent advent of large language models has catalyzed the development of vision-language Models (VLMs) (Li et al., 2024c; Wang et al., 2024; Achiam et al., 2023; Team et al., 2023; Deepmind, [n. d.]), which offer a unified natural language interface for diverse vision tasks and present a promising avenue for analyzing user videos real time. Despite their potential, current VLMs face substantial hurdles when used for physical task guidance, especially for BLV users. One primary issue is hallucination (Bai et al., 2024), where models produce information not grounded in the visual evidence. Also, these models often generate guidance presupposing visual capabilities, leading to suggestions unsuitable for BLV users (Bigham et al., 2010; Song et al., 2024). Further, existing VLMs often lack fine-grained analysis capabilities (Xue et al., 2024a), struggling to monitor detailed execution progress within tasks, which is vital for step-by-step assistance. In this work, we carefully design task support grounded in BLV people’s needs in real-world scenarios and propose a pipeline to generate accessible task guidance.
@@ 3.1. Method
Figure 3. Timelines of utterances annotation of observational study sessions. Colors denote utterance types and the x-axis plots interaction order. VRT2 provided more proactive supplementary information and progress descriptions to novice BLV2.
@@ 3.3. Study Findings
Takeaway: Provide detailed instructions based on both the narration and visuals from the video. Structure rich information into clear, concise segments to support easier understanding.
@@
Takeaway: Supplement instructions with tips and accessible workarounds for both immediate guidance and building reusable strategies for future tasks.
@@
Takeaway: Provide proactive visual feedback on user progress and success, and prompt users when adjustments are needed to keep relevant actions in view.
@@
Takeaway: Elicit users to leverage non-visual sensory cues to evaluate the progress, to build confidence and support independent task completion.
@@
Takeaway: Address users’ diverse questions by grounding responses in both the user’s current task progress and the how-to video knowledge, enabling flexible task execution.
@@
Takeaway: Adapt guidance to user skills, preferences, and context by adjusting the level of detail, rephrasing based on familiarity, and referencing available tools or spatial layout.
@@ 5.2. Supplementing Accessible Strategies
Retrieval Augmented Generation. General-purpose AI suggestions often assume dominant user groups, and may generate irrelevant suggestions that make BLV users feel misfit (e.g., Check if the butter looks golden brown(Alharbi et al., 2024). Even when prompted to provide accessible tips, these models reflect biased assumptions and suggest that users seek help from sighted people or offer overly generic advice without practical value – reminding them to be careful while cooking. To address these limitations, we use a multi-modal retrieval augmented generation (RAG) approach (Lewis et al., 2020) to ground AI suggestions in real-world, disability-aware resources. From the curated dataset, our pipeline first extracts text chunks to generate text embeddings, and identifies images and describes them with a VLM (Gemini 1.5 pro) to generate embeddings. At inference time, we input queries to generate step-specific workarounds, and the query also incorporates user-specific context to further personalize suggestions (e.g., vision level, cooking experience, available tools and kitchen setup) (see § 8 for details). We compute the cosine similarity between the query’s text embedding and the stored embeddings and retrieve top-3 text chunks to generate answers with the context. It then aggregates these text chunks into a tip, surfacing multiple alternatives when appropriate (Figure 1.2).
@@ 5.3. Progress Monitoring
From the how-to video, our pipeline generates a set of criteria (irrelevant, in-progerss, complete, mistake) for each action and uses these to monitor the user’s progress (Figure 5). These criteria are derived using a VLM (Achiam et al., 2023) (details in §C). Instead of directly comparing frames from the how-to and user streams, we use abstracted criteria that are more robust to variations in tools, occlusions, and user behavior. As users perform each action, Vid2Coach uses these criteria to adapt its feedback: it pauses during unrelated or impromptu actions (e.g., washing hands), provides progress updates while a step is in-progress, and proactively suggests moving on once the step appears complete.  Prior work demonstrates that people adapt recipes by combining, skipping, or improvising steps (Ashutosh et al., 2024b; Yang et al., 2025; Comber et al., 2013), a pattern mirrored by the VRTs we observed. Thus, Vid2Coach lets users freely navigate the steps and evaluates each step independently rather than assuming a strict step order (Li et al., 2025a).
To make this system effective in real-time streamed video, we propose a 3-way categorization of actions – punctual, iterative, or durative – based on its temporal characteristics. Punctual actions (e.g.,“add 1 cup of flour”) happen quickly and may be missed in frame sampling or batched inference. For punctual actions, in-progress feedback is often untimely or unnecessary, so Vid2Coach only confirms completion. For iterative actions (e.g., “place three scoops of cookie dough”) that involve repetition, it counts visible repetitions for tracking progress. For durative actions (e.g., “cook until golden brown”) that involve gradual visual change, Vid2Coach offers real-time updates on its progression and completion. For actions where visual cues are minimal (e.g., “wait for the mixture to cool down”), Vid2Coach asks users to decide when to proceed. Vid2Coach prevents false positive step progression by prompting a quick confirmation: – “You seem to be complete because the butter looks golden brown. Would you like to move on to the next step?”
To balance the accuracy and speed of realtime feedback, we adopt a dual-model approach. Inspired by WorldScribe (Chang et al., 2024), which leveraged different models with varying latency and granularity, our setup combines 1) a batch model (Gemini 2.0) for deeper reasoning with 2) a lightweight streaming model (Gemini 2.0-Live) for immediate responsiveness. The batch VLM runs every 5 seconds and looks at the five most recent frames (sampled at 1 fps) to determine user status based on pre-generated completion criteria. Unlike batch models, streaming models begin generating output token-by-token before fully processing the input (Xiao et al., 2023), making them less suited for complex reasoning but valuable for continuous monitoring. Thus, we use the streaming VLM to provide real-time, low-latency descriptions of user actions when a durative action is in progress. Users can toggle this progress feedback on or off using a voice command. To avoid memory buildup and maintain inference consistency, Vid2Coach resets the model’s input context at each step, preventing cascading errors in the pipeline.
@@ 5.4. Intent Elicitation & Question Answering
In addition to automated feedback, Vid2Coach also responds to users’ open-ended, voice-based queries. To ensure responsiveness, we use the streaming model’s voice activity detection (VAD) to immediately interrupt any ongoing generation and prioritize the user’s spoken query. Inspired by patterns observed in our formative study (§3), we implement a few-shot classification prompt that identifies the user’s intent from both explicit utterances (e.g., “Does this look complete?”) and implicit expressions (e.g., “I can’t tell which of these is sugar”). We classify user utterances into the five intent types and trigger a different function call to handle each request:
@@ 6.1. Generating Instructions
Results. As table 1 summarizes, Vid2Coach outperformed baselines by identifying more task-relevant atomic facts and omitting fewer details present in the original narration. As shown in Figure 6, Vid2Coach surfaced concrete visual details that can support BLV users to accurately follow the steps (e.g., describing the use of a measuring spoon, amount of bacon used, and how they’re presented). These details are especially important for BLV users, who might otherwise miss visual cues and inaccurately follow the action.
@@ 7.4. Limitations and Opportunities
Future Opportunities. Participants often adapted the step sequence by reordering steps for parallel task execution or skipping steps based on dietary preferences. However, this flexibility sometimes led to unintentional skips. In the future, intent can be inferred by analyzing interaction patterns such as rapid, sequential skips (often errors) vs. deliberate reordering, and checking whether the new action sequence still results in a valid completion of the recipe (Idrees et al., 2023). Also, Vid2Coach can mark core steps whose omission would compromise safety or outcome (e.g., turning off a burner) and optional steps (e.g., sprinkling cinnamon), and confirm users for skipping core steps. As steps are interdependent, skipping or altering one affect subsequent ones (e.g., leaving bread unsliced changes cooking time). We can leverage step dependencies for flexible guidance.
Participants also saw Vid2Coach’s potential to enable independent task skill learning in everyday lives. Compared to human assistants, participants found Vid2Coach more readily available, task-relevant, and nonjudgmental. Several participants emphasized the personal motivation and sense of achievement they felt by completing tasks on their own, even when human help was accessible. P4 noted “Even if sighted volunteers are available 24/7 for free, I want to know I can achieve it on my own.” P11, who saw the potential in personalization of Vid2Coach said “Unlike human assistants whose description styles change every call, I can train it [Vid2Coach] to be personalized to me.” P5 also reflected on societal opportunities “I think it can help blind students get more jobs. A lot of sighted people assume blind people can’t learn new skills quickly or follow manuals. But with this, we can.”
@@ 9. Discussion
9.1. Video as a Knowledge Source
@@ 9.3. Expanding to New Activity Domains
9.4. Scaling Instructional Knowledge
While Vid2Coach uses curated accessibility guidelines and BLV-shared practices for common tasks like cooking and crafting, resources for less common tasks (e.g., furniture assembly, planting) are limited. Future systems can incorporate community-sourced knowledge, while also actively eliciting and supporting the sharing of such knowledge. This could involve creating lightweight platforms for BLV users to exchange non-visual workarounds, similar to how blind communities share tips through forums and mailing lists. Peer-to-peer knowledge exchange may surface effective strategies that may not yet be captured in formal accessibility resources. To reduce the time and cost for dataset collection, future work can consider LLM-based agents that automatically crawl task-relevant documents and verify credibility (Huang et al., 2024). Vid2Coach can also refine or expand retrieval queries to better handle steps unseen in the dataset. For example, it can broaden the search by transforming a specific and complicated query into generic sub-queries (Chan et al., 2024), or substituting terms with similar ingredients/tools (Lei et al., 2024). For domains like exercise, dance, and the arts, generic written tips are scarce and often less useful than personalized actionable feedback, as the task success depends more on body control and technique. To detect deviations from a reference action and generate concrete correction feedback (e.g., lift your elbow slightly higher), future systems can analyze users’ live hand- or body-pose data or leverage expert feedback dataset (Ashutosh et al., 2024a). Finally, hybrid models can offer fallback support with human experts (VRTs) and peer-support communities, and learn human responses and automate itself over time (Huang et al., 2018).
9.5. Beyond Voice Assistants
@@
9.6. Complementary Expert-AI Support
The field of Virtual Reality Therapy (VRT) is relatively small and growing more slowly than the increasing demand (Connors et al., 2023). As VRTs teach a wide range of life skills (e.g., navigation, technology), most of their time goes to foundational, in-person training, leaving little capacity for personalized recipe coaching (Section 3.3). Vid2Coach addresses this gap by supporting BLV users in independently learning new recipes after they have received foundational training (e.g., holding a knife) from VRTs. VRTs in our formative study were excited for such systems, recognizing their potential to extend support between in-person sessions. Future HCI work can explore how VRTs and systems like Vid2Coach can collaborate more effectively, supporting long-term training programs. Just as in physical therapy or music lessons — where basics are taught in person and home exercises are assigned (Argent et al., 2018) — future VRTs can give basic training, assign recipe videos matched to students’ skill level, and review the user’s cooking session to offer nuanced coaching and encouragement. Systems can flag challenge points (e.g., moments of hesitation or unsafe motion) (Mysore and Guo, 2018) and let the VRT select or record brief corrective clips that the system bundles into a learner-specific practice set. In this workflow, the system handles detection and summarization, while the VRT provides the expert guidance, blending automated detection with human expertise in complementary roles.